OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4

OpenStudy (anonymous):

3(x+1)=-2(x-1)-4 distributive property 3x+3=-2x-2+4 add 2x to both sides 5x-3=2 add three to both sides 5x=5 divide 5 by both sides x=1